OpenStudy (anonymous):

What values for theta (0<=theta<=2pi) satisfy the equation 2 cos theta+1=-cos theta?

OpenStudy (anonymous):

\[ \theta (0\le \theta \le 2 \pi ) \\ 2 \cos \theta+1=-\cos \theta\]

OpenStudy (jdoe0001):

\(\bf 2 cos (\theta)+1=-cos(\theta) \ \ ?\)

OpenStudy (anonymous):

here are the choices

OpenStudy (anonymous):

i don't really know where to start on this one at all

OpenStudy (jdoe0001):

is it => \(\bf \bf 2 cos (\theta+1)=-cos(\theta) \ \ ?\) or \(\bf \bf 2 cos (\theta)+1=-cos(\theta) \ \ ?\)

OpenStudy (anonymous):

no its the second

OpenStudy (anonymous):

or wait no i think you're right i misread that

OpenStudy (anonymous):

it is 2 cos theta+1

OpenStudy (jdoe0001):

hmmm no, I think you're right, is \(\bf 2 cos (\theta)+1=-cos(\theta)\)

OpenStudy (jdoe0001):

\(\bf 2 cos (\theta)+1=-cos(\theta) \implies 3 cos (\theta)+1 = 0 \implies cos (\theta) = \\ cos^{-1}(cos (\theta)) = cos^{-1}\left(-\cfrac{1}{3}\right) \implies \theta = cos^{-1}\left(-\cfrac{1}{3}\right)\) then just get that, arcCosine in Radians, as far as I can tell is the choices

OpenStudy (jdoe0001):

hmmm

OpenStudy (jdoe0001):

\(\bf 2 cos (\theta)+1=-cos(\theta) \implies 3 cos (\theta)+1 = 0 \implies cos (\theta) = -\cfrac{1}{3}\\ cos^{-1}(cos (\theta)) = cos^{-1}\left(-\cfrac{1}{3}\right) \implies \theta = cos^{-1}\left(-\cfrac{1}{3}\right)\)

OpenStudy (anonymous):

so is it 1.91 and 4.37?

OpenStudy (jdoe0001):

well, what did you get?

OpenStudy (jdoe0001):

ohh, I see, you got 1.91

OpenStudy (anonymous):

is that right

OpenStudy (jdoe0001):

yes, that's it, see the interval is \(\bf \theta \in (0\le \theta \le 2 \pi )\) 0 = 0 Radians \(\bf 2\pi = 2\times 3.1416 \implies 6.28\) so 1.91 and 4.37 Radians are well within that rage

OpenStudy (jdoe0001):

range rather

OpenStudy (anonymous):

yay thank you, i usually always guess but now I kind of understand